# EvAlignUX: Advancing UX Research through LLM-Supported Exploration of Evaluation Metrics

### v1

@@ Abstract.
Evaluating UX in the context of AI’s complexity, unpredictability, and generative nature presents unique challenges. HCI scholars lack sufficient tool support to build knowledge around diverse evaluation metrics and develop comprehensive UX evaluation plans. In this paper, we introduce EvAlignUX , an innovative system grounded in scientific literature and powered by large language models (LLMs), designed to help HCI scholars explore evaluation metrics and their relationship to potential research outcomes. A user study involving 19 HCI scholars revealed that EvAlignUX significantly improved the perceived clarity, specificity, feasibility, and overall quality of their evaluation proposals. The use of EvAlignUX enhanced participants’ thought processes, resulting in the creation of a Question Bank that can be used to guide UX Evaluation Development. Additionally, the influence of researchers’ backgrounds on their perceived inspiration and concerns about over-reliance on AI highlights future research directions for AI’s role in fostering critical thinking.
@@ 1. Introduction
The rapid integration of AI into interactive systems presents a growing challenge for user experience (UX) evaluation (Stige et al., 2023; Chaudhry, 2024). Traditional usability evaluation guided by standard frameworks such as the System Usability Scale (SUS) (Bangor et al., 2008), User Experience Questionnaire (UEQ) (Schrepp et al., 2017), NASA Task Load Index (NASA-TLX) (Hart, 2006), fall short in capturing the complex, dynamic interaction between humans and AI (Kuang et al., 2024). As AI systems become more context-aware and capable of generating responses in real-time, UX researchers are confronted with new questions about trust, emotional engagement, and ethical considerations (Liao et al., 2020a; Ahmad et al., 2023a; Shneiderman, 2020; Ntoa et al., 2021). Addressing these new considerations is especially challenging for systems research, where a range of evaluation methods, often borrowing and combining techniques from software engineering, design, and usability evaluation, is triangulated to evaluate the UX of interactive computing systems (Marquardt et al., 2017). Current UX evaluation methods, e.g., Post-Study System Usability Questionnaire (PSSUQ) (Lewis, 1992), Subjective Mental Effort Questionnaire (SMEQ) (Hyland et al., 1989), while robust for traditional systems, are inadequate for AI-powered interfaces. Researchers are striving to develop new frameworks, which take into account ethical considerations, trust dynamics, intimacy, and multimodal behaviors, to adequately guide scholars and practitioners in developing their UX evaluation plans for human-AI interaction research, e.g., (Zheng et al., 2022; Liao et al., 2024). Given the plethora of well documented evaluation methods, heuristics and metrics (e.g., (Lazar et al., 2017; Nielsen, 1992; Vermeeren et al., 2010)), much less is known about How researchers select appropriate metrics to conduct a comprehensive evaluation of a system.
@@
Also, despite the availability of various UX evaluation methods (Pettersson et al., 2018; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019; Kieffer et al., 2019), there is a lack of practical tools to guide researchers in developing comprehensive evaluation plans. Research has shown that incorporating evaluation metrics as part of the design materials in early stages can improve the clarity of designing human-AI interaction research (Zheng and Huang, 2023). UX is frequently regarded as a rigid set of methods, leading to calls for research and education that encourage framing UX as “more of a mindset than a method” within the community (Gray, 2016). However, with the rapid advancement of AI models, the literature on AI-mediated multi-stakeholder interactions has expanded exponentially. This growth presents challenges for novice researchers in building up knowledge of relevant evaluation metrics, while seasoned researchers also face increasing pressure to stay current with the latest developments.
@@
To bridge this gap, we designed, implemented, and evaluated EvAlignUX , an interactive system that supports researchers in developing UX evaluation plans for their research projects. EvAlignUX is powered by the generative capabilities of Large Language Models (LLMs) to facilitate the exploration of UX evaluation ideas, as LLMs have demonstrated their potential for ideation of scientific research by processing a vast amount of information, generating context-aware insights, retrieving literature, and brainstorming (Doshi and Hauser, 2024; Glickman and Zhang, 2024; Liu et al., 2024). Specifically, as shown in Fig. 1, EvAlignUX comprises three primary panels: the Project Ideation Panel, the Metric Explorer Panel, and the Outcomes and Risks Panel. These panels allow users to explore UX metrics, potential outcomes, and potential risks associated with human-AI interaction research. In an evaluation with 19 HCI scholars, EvAlignUX led to significant improvements in the clarity, concreteness, and confidence in their UX proposals. The results demonstrated that EvAlignUX significantly improved the clarity, concreteness, and confidence of researchers in their proposed UX evaluation plans. Participants were able to identify and integrate detailed UX metrics, self-critique their plans, and raise awareness of potential risks using the system features.
@@
This work makes several significant contributions to the HCI community. First, EvAlignUX is a novel system, designed to support the ideation process of developing UX evaluation plans. Powered by LLMs, EvAlignUX enables users to explore UX metrics and their relationships with potential research outcomes. The unique visual interface of the metrics graph supports users in examining the connections between different UX metrics. To the best of our knowledge, EvAlignUX fills the gap in tool support for researchers, providing insights into the research literature of human-AI evaluation metrics and facilitating better decision-making regarding UX evaluation. Second, our user studies demonstrate the effectiveness of EvAlignUX in helping users gain significantly more clarity and feasibility in their proposed UX evaluation plans. Our findings revealed that researchers’ background impacted their perceived levels of inspiration and their concerns about over-reliance on AI. They also integrate the system features into their thought processes differently, focusing on varied utilities such as comprehensiveness of self-critique and efficiency of implementing the ideas. Third, our analysis of participants’ think-aloud and perceived value of different features produced a comprehensive Question Bank, highlighting the specific questions researchers asked at different steps of UX evaluation development and identifying how AI provide valuable support such as envisioning the research impact, assessing the risks involved in the research, and justifying metric selection. Lastly, our findings prompt a timely and profound discussion on shifting the focus from method to mindset in future research on UX evaluation.
@@ 2. Related Work
In this section, we review the literature on (1) current UX evaluation approaches, and (2) the potential of leveraging LLMs to provide tool support.
@@ 2.1. UX Evaluation of HCI System Research
Developing a UX evaluation plan can be challenging, as it’s often unclear what constitutes ”evaluation” and which metrics are appropriate (Ledo et al., 2018; Hillman et al., 2023). Hillman et al. (Hillman et al., 2023) elaborated on several challenges involved in developing a solid evaluation plan, such as selecting metrics that accurately reflect UX and research goals, choosing data collection methods, establishing collection protocols, and ensuring consistency across products and organizations. Additionally, long-term effects and social or cultural barriers further complicate the process (Hillman et al., 2023; Darin et al., 2019; Vermeeren et al., 2010). In the past, HCI scholars often relied on established surveys like System Usability Scale (SUS), User Experience Questionnaire (UEQ), Post-Study System Usability Questionnaire (PSSUQ), NASA Task Load Index (NASA-TLX), Subjective Mental Effort Questionnaire (SMEQ) and Customer Satisfaction (CSAT) to measure users’ perceptions (Bangor et al., 2008; Schrepp et al., 2017; Lewis, 1992; Hart, 2006; Reichheld, 2003). PULSE metrics (Page views, Uptime, Latency, etc.) are also commonly used to track user behavior (Rodden et al., 2010). However, these metrics are often “low-level or indirect measures of user experience” (Rodden et al., 2010). Several evaluation frameworks relating to user-centered metrics have been introduced to the HCI and UX domains, such as the Software Development Productivity Framework (Sadowski et al., 2019), and Google’s HEART framework (Rodden et al., 2010). For instance, the HEART framework, which includes Happiness, Engagement, Adoption, Retention, and Task success, was introduced for measuring web-based applications at scale (Rodden et al., 2010).
@@
With the rise of AI, recent studies on UX research lay out several new trends and challenges. A school of HCI researchers aims to identify and propose new evaluation metrics for both usability and user-centered evaluation in the age of AI. For example, Zheng et al.’s literature review of human-AI interaction research in HCI, which included user evaluations, reported a suite of metrics using a survey, system logs, and interviews to collect (Zheng et al., 2022) and argued the importance of including more user-centered metrics (e.g., sociability, socio-emotional, and AI anthropomorphism metrics) with AI’s evolving role in human social interactions. Likewise, Faruk et al.’s works also proposed new subjective scales for measuring the UX of voice agents (Faruk et al., 2024a) and conversational agents (Faruk et al., 2024b). In other domains, like healthcare, Abbasian et al. also summarized some foundation metrics for evaluating the effectiveness of healthcare conversations powered by generative AI (Abbasian et al., 2024).
@@
However, there is a lack of tools to help researchers understand the evaluation metrics literature and design a comprehensive evaluation plan for human-AI interaction research. As Yang et al. (Yang et al., 2020) systematically reported how AI system’s uncertainty and complexity can make unpredictable errors, which easily damage UX and lead to undesired societal impact. Gray’s finding of ”More of a Mindset Than a Method” also highlighted that UX practitioners often follow a codified method rather than adopting a ”mindset” that aligns with their design process, practice context, and the specific design problem (Gray, 2016). All these works suggest that new tools and approaches are needed to support researchers and practitioners in developing their UX evaluation plans.
2.2. The Potential of LLMs to Support Ideation
@@
Large Language Models (LLMs) can process vast amounts of multi-modal information and generate context-aware, coherent text, making them ideal tools for evaluation-related tasks in various domains. A group of researchers developed novel interaction prototypes using LLMs for evaluating generative outputs (Desmond et al., 2024; Kim et al., 2024; Shankar et al., 2024). For example, Kim et al. developed EvalLM (Kim et al., 2024), an interactive system for refining prompts by evaluating multiple outputs on user-defined criteria. This helps developers compare their prompt outputs to diagnose weaknesses in their prompts. Regarding LLMs’ support for scientific tasks, recent studies have showcased LLM-powered methods for retrieving and recommending relevant literature (Shen et al., 2023), reviewing and critiquing research drafts, writing research manuscripts, and augmenting research ideation and brainstorming for scientific writing based on scientific evidence and database (Nigam et al., 2024), utilizing large scientific dataset. Recent advancements in HCI and AI research also attempt to support UX practitioners with AI-enabled tools to conduct usability testing (Fan et al., 2022; Kuang et al., 2024), e.g., LLMs are employed to help human evaluators to identify more usability issues in usability test video recordings. All these research highlights the significant potential of LLMs to support research ideation and provide valuable feedback for task evaluation. However, little is known about how researchers with varying levels of experiences can benefit from using LLM-supported tools to develop their UX evaluation plans.
@@ 3. EvAlignUX
To address the research gap, EvAlignUX was designed as a tool to aid researchers explore evaluation metrics for developing UX plans of Human-AI research. The system was developed through an iterative process, informed by both the literature review (Zheng et al., 2022; Inan Nur et al., 2021; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019) and input from a focus group. Leveraging the authors’ expertise in HCI research and practices, we began with an initial process of developing UX evaluation plan, e.g., proposing of a research topic, identifying relevant UX metrics for evaluation, and then framing the expected research contribution. The design was refined over two months through multiple weekly feedback sessions with an HCI research group, consisting of nine graduate students and faculty members from the HCI, Informatics, and Computer Science departments (n=9). All of the participants have worked on at least one HCI project where they designed or developed systems. Their applications covered diverse contexts: vision-based question-answering, ideation and writing, accessibility, immersive technologies, and intelligent assistants. Based on the UX evaluation challenges reported from prior literature (Ledo et al., 2018; Hillman et al., 2023), and after piloting the initial pipeline with this group, we established three design goals (DG), then we iteratively searched for prior literature and consulted for feedback through this weekly focus group meetings.
DG1: Recommend UX metrics relevant for a project;
DG2: Explain how the metrics are used in existing literature; and
DG3: Inform researchers with potential risks associated with their design and evaluation plans.

### v2

# EvAlignUX: Advancing UX Research through LLM-Supported Evaluation Metrics Exploration

@@ Abstract.
Evaluating UX in the context of AI’s complexity, unpredictability, and generative nature presents unique challenges. How can we support HCI researchers to create comprehensive UX evaluation plans? In this paper, we introduce EvAlignUX , a system powered by large language models and grounded in scientific literature, designed to help HCI researchers explore evaluation metrics and their relationship to research outcomes. A user study with 19 HCI scholars showed that EvAlignUX improved the perceived quality and confidence in UX evaluation plans while prompting deeper consideration of research impact and risks. The system enhanced participants’ thought processes, leading to the creation of a “UX Question Bank” to guide UX evaluation development. Findings also highlight how researchers’ backgrounds influence their inspiration and concerns about AI over-reliance, pointing to future research on AI’s role in fostering critical thinking. In a world where experience defines impact, we discuss the importance of shifting UX evaluation from a “method-centric” to a “mindset-centric” approach as the key to meaningful and lasting design evaluation.
The image presents the interface of EValignUX, divided into three main panels: Ideation Panel (left), Metrics Explorer Panel (center), and Impact and Risk Panel (right). The Ideation Panel contains a user-provided system description (1) and an initial UX evaluation plan (2), along with an index field (A) that suggests relevant indexes, like ‘application domain’ and ‘paradigm’ based on the project description. The Metrics Explorer Panel features a metrics graph (B) showing connections between UX metrics, which visualizes features as nodes with varying sizes and colors to indicate their importance. Below the graph is a metrics list (C), offering detailed descriptions of suggested metrics like ‘Attitude Towards the Agent’ and ‘Closeness’. The Impact and Risk Panel allows users to define expected UX evaluation outcomes (3) like ‘an accessible, affordable, and personalized online counseling chatbot’ (D), while also listing potential risks (E) such as ‘Unauthorized Data’ and ‘Incomplete Data Attribute Capture’.
@@ 1. Introduction
The rapid integration of AI into interactive systems presents a growing challenge for user experience (UX) evaluation (Stige et al., 2023; Chaudhry, 2024). Traditional usability evaluation guided by standard frameworks such as the System Usability Scale (SUS) (Bangor et al., 2008), User Experience Questionnaire (UEQ) (Schrepp et al., 2017), NASA Task Load Index (NASA-TLX) (Hart, 2006), falls short in capturing the complex, dynamic interaction between humans and AI (Kuang et al., 2024). As AI systems become more context-aware and capable of generating responses in real-time, UX researchers are confronted with new questions about trust, emotional engagement, and ethical considerations (Liao et al., 2020a; Ahmad et al., 2023a; Shneiderman, 2020; Ntoa et al., 2021).
Addressing these new considerations is especially challenging for systems research, where a range of evaluation methods, often borrowing and combining techniques from software engineering, design, and usability evaluation, is triangulated to evaluate the UX of interactive computing systems (Marquardt et al., 2017). Current UX evaluation methods, e.g., the Post-Study System Usability Questionnaire (PSSUQ) (Lewis, 1992), the Subjective Mental Effort Questionnaire (SMEQ) (Hyland et al., 1989), while robust for traditional systems, are inadequate for AI-powered interfaces. Researchers are striving to develop new frameworks that take into account ethical considerations, trust dynamics, intimacy, and multimodal behaviors, to adequately guide scholars and practitioners in developing their UX evaluation plans for human-AI interaction research, e.g., (Zheng et al., 2022; Liao et al., 2024). Given the plethora of well-documented evaluation methods, heuristics, and metrics (e.g. (Lazar et al., 2017; Nielsen, 1992; Vermeeren et al., 2010)), much less is known about How researchers select appropriate metrics to conduct a comprehensive evaluation of a system.
@@
Also, despite the availability of various UX evaluation methods (Pettersson et al., 2018; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019; Kieffer et al., 2019), there is a lack of practical tools to guide researchers in developing comprehensive evaluation plans. Research has shown that incorporating evaluation metrics as part of the design materials in the early stages can improve the clarity in the design of human-AI interaction research (Zheng and Huang, 2023). UX is frequently regarded as a rigid set of methods, leading to calls for research and education that encourage framing UX as “more of a mindset than a method” within the community (Gray, 2016). However, with the rapid advancement of AI models, the literature on AI-mediated multi-stakeholder interactions has expanded exponentially. This growth presents challenges for novice researchers in acquiring knowledge of relevant evaluation metrics, while seasoned researchers also face increasing pressure to stay current with the latest developments.
@@
To bridge this gap, we designed, implemented, and evaluated EvAlignUX , an interactive system that supports novice and experienced HCI researchers in developing UX evaluation plans for their research projects. EvAlignUX is powered by the generative capabilities of Large Language Models (LLMs) to facilitate the exploration of UX evaluation ideas, as LLMs have demonstrated their potential for the ideation of scientific research by processing vast amount of information, generating context-aware insights, retrieving literature, and brainstorming (Doshi and Hauser, 2024; Glickman and Zhang, 2024; Liu et al., 2024). Specifically, as shown in Fig. 1, EvAlignUX comprises three primary panels: the Project Ideation Panel, the Metric Explorer Panel, and the Outcomes and Risks Panel. These panels allow users to explore UX metrics, potential outcomes, and potential risks associated with human-AI interaction research. In an evaluation with 19 HCI scholars, EvAlignUX led to significant improvements in the clarity, concreteness, and confidence of their revised UX evaluation plans. Participants were able to better justify selected UX metrics, self-critique their plans, and raise awareness of potential risks using the system features.
@@
This work makes several significant contributions to the HCI community. First, EvAlignUX is a novel system designed to support the ideation process of developing UX evaluation plans. Powered by LLMs, it enables users to explore UX metrics and their relationships with potential research outcomes. The unique visual interface of the metrics graph supports users in examining the connections between different UX metrics. To the best of our knowledge, EvAlignUX fills the gap in tool support available to researchers by providing insights into the research literature on human-AI evaluation metrics and facilitating better decision-making regarding UX evaluation. Second, our user studies demonstrate the effectiveness of EvAlignUX in helping users achieve significantly greater clarity and feasibility in their proposed UX evaluation plans. Our findings revealed that researchers’ backgrounds impacted their perceived levels of inspiration and their concerns about over-reliance on AI. They also integrated the system features into their thought processes differently, focusing on varied utilities such as the comprehensiveness of self-critique and the efficiency of implementing the ideas. Third, our analysis of participants’ think-aloud protocols and perceived value of different features produced a comprehensive Question Bank for UX Evaluation Development, highlighting the specific questions researchers asked at different steps of UX evaluation development and identifying how AI provides valuable support such as envisioning the research impact, assessing the risks, and justifying metric selection. Lastly, our findings prompt a critical discussion on shifting the focus from method to mindset in future research on UX evaluation.
@@ 2. Related Work
In this section, we review the literature on (1) current UX evaluation approaches and (2) the potential of leveraging LLMs to provide tool support.
@@ 2.1. UX Evaluation of HCI System Research
Developing a UX evaluation plan can be challenging, as it is often unclear what constitutes ”evaluation” and which metrics are appropriate (Ledo et al., 2018; Hillman et al., 2023). Hillman et al. (Hillman et al., 2023) elaborated on several challenges involved in developing a solid evaluation plan, such as selecting metrics that accurately reflect UX and research goals, choosing data collection methods, establishing collection protocols, and ensuring consistency across products and organizations. Additionally, long-term effects and social or cultural barriers further complicate the process (Hillman et al., 2023; Darin et al., 2019; Vermeeren et al., 2010). Several foundational conceptual frameworks serve as the backbone for UX evaluation, including Nielsen’s usability engineering (Nielsen, 1994b) and heuristic evaluation (Nielsen, 1994a), Hassenzahl’s UX framework (Hassenzahl et al., 2010), and Norman’s emotional design theory (Norman, 2007). These frameworks emphasize multi-dimensional assessments, moving beyond performance metrics to encompass user perceptions, attitudes, and context of use.
@@
To measure users’ perceptions quantitatively, researchers often use established surveys, such as the System Usability Scale (SUS), the User Experience Questionnaire (UEQ), the Post-Study System Usability Questionnaire (PSSUQ), the NASA Task Load Index (NASA-TLX), the Subjective Mental Effort Questionnaire (SMEQ), and the Customer Satisfaction (CSAT) (Bangor et al., 2008; Schrepp et al., 2017; Lewis, 1992; Hart, 2006; Reichheld, 2003). PULSE metrics (e.g., page views, uptime, latency, etc.) are also commonly used to track user behavior (Rodden et al., 2010). However, these metrics are often “low-level or indirect measures of user experience” (Rodden et al., 2010). Several evaluation frameworks relating to user-centered metrics have been introduced to the HCI and UX domains, such as the Software Development Productivity Framework (Sadowski et al., 2019) and Google’s HEART framework (Rodden et al., 2010). For instance, the HEART framework, which includes Happiness, Engagement, Adoption, Retention, and Task Success, was introduced to measure web-based applications at scale (Rodden et al., 2010).
Beyond selecting appropriate metrics to evaluate systems, researchers have also developed methods and tools to enhance survey and interview question design. For example, the Question Appraisal System (QAS-99) (Willis and Lessler, 2013) and Willis’s cognitive interviewing methodology (Willis, 2004) provide structured approaches for detecting and refining data collection items, ensuring that questions are clear, unbiased, and aligned with research goals. Similarly, QUAID (Graesser et al., 2006) offers diagnostic support for identifying comprehension issues in survey questions, while Survey Sidekick (Hsiao et al., 2014) assists in creating scientifically sound surveys. Semi-automated approaches (Siemon et al., 2016) further aid in generating relevant and effective questions. Incorporating such methods into UX evaluation can yield instruments that better capture users’ experiences, perceptions, and emotional responses.
@@
As HCI rapidly shifts toward AI-driven systems, recent studies in UX research highlight several emerging trends and challenges. A group of HCI researchers aims to identify and propose new evaluation metrics for both usability and user-centered evaluation in the age of AI. For example, Zheng et al.’s literature review of human-AI interaction research in HCI, which included user evaluations, reported a suite of metrics that were collected using surveys, system logs, and interviews to collect (Zheng et al., 2022) and argued the importance of including more user-centered metrics (e.g., sociability, socio-emotional, and AI anthropomorphism metrics) with AI’s evolving role in human social interactions. Likewise, Faruk et al. proposed new subjective scales for measuring the UX of voice agents (Faruk et al., 2024a) and conversational agents (Faruk et al., 2024b). In other domains, such as healthcare, Abbasian et al. summarized some foundation metrics for evaluating the effectiveness of healthcare conversations powered by AI (Abbasian et al., 2024).
@@
However, there is a lack of tools to help researchers understand the evaluation metrics literature and design a comprehensive evaluation plan for human-AI interaction research. As Yang et al. (Yang et al., 2020) systematically reported that an AI system’s uncertainty and complexity can lead to unpredictable errors that damage UX and result in undesired societal impact.. Gray’s finding, encapsulated in the phrase ”more of a mindset than a method” also highlighted that UX practitioners often follow a codified method rather than adopting a ”mindset” that aligns with their design process, practice context, and the specific design problem (Gray, 2016). All these works suggest that new tools and approaches are needed to support researchers and practitioners in developing their UX evaluation plans.
2.2. The Potential of LLM-Supported Ideation
@@
Large Language Models (LLMs) can process vast amounts of multi-modal information and generate context-aware, coherent text, making them ideal tools for evaluation-related tasks in various domains. A group of researchers developed novel interaction prototypes using LLMs for evaluating generative outputs (Desmond et al., 2024; Kim et al., 2024; Shankar et al., 2024). For example, Kim et al. developed EvalLM (Kim et al., 2024), an interactive system for refining prompts by evaluating multiple outputs on user-defined criteria. This helps developers compare prompt outputs to diagnose weaknesses in their prompts. Regarding LLMs’ support for scientific tasks, recent studies have showcased LLM-powered methods for retrieving and recommending relevant literature (Shen et al., 2023), reviewing and critiquing research drafts, writing research manuscripts, and augmenting research ideation and brainstorming for scientific writing by leveraging scientific evidence and large datasets (Nigam et al., 2024). Recent advancements in HCI and AI research also have also aimed to support UX practitioners with AI-enabled tools for usability testing (e.g., LLMs are employed to help human evaluators identify additional usability issues in usability test video recordings). All these research highlights the significant potential of LLMs to support research ideation and provide valuable feedback for task evaluation. However, little is known about how researchers with varying levels of experience can benefit from using LLM-supported tools to develop their UX evaluation plans.
@@ 3. EvAlignUX
To address the research gap, EvAlignUX was designed as a tool to aid researchers in exploring evaluation metrics for developing UX plans for Human-AI research. The system was developed through an iterative process, informed by both the literature review (Zheng et al., 2022; Inan Nur et al., 2021; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019) and input from a focus group. Leveraging the authors’ expertise in HCI research and practices, we began with an initial process of developing a UX evaluation plan, e.g., proposing a research topic, identifying relevant UX metrics for evaluation, and framing the expected research contribution. The design was refined over two months through multiple weekly feedback sessions with an HCI research group composed of, consisting of nine graduate students and faculty members from the HCI, Informatics, and Computer Science departments (n=9). All of the participants have worked on at least one HCI project where they designed or developed systems. Their applications covered diverse contexts: vision-based question-answering, ideation and writing, accessibility, immersive technologies, and intelligent assistants. Based on the UX evaluation challenges reported in prior literature (Ledo et al., 2018; Hillman et al., 2023), and after piloting the initial pipeline with this group, we established three design goals (DG). We then iteratively searched for additional literature and sought feedback through weekly focus group meetings.
DG1: Recommend UX metrics relevant to a project;
DG2: Explain how the metrics are used in existing literature;
DG3: Inform researchers about potential risks associated with their design and evaluation plans.